Business and Finance. When x number of units are sold, the price of each unit (in dollars) is given by p= -x/2+75. Find the unit price when the following quantities are sold: 2,7,9,11

-x/2 is a fraction.
Please help me.

Answer by tutorcecilia(2152)

Plug-in each value of x (2,7,9,11)into the equation (p=-x/2+75) and solve for x:
p= -(2)/2+75=74
p= -(7)/2+75=71.50
p= -(9)/2+75=70.50
p= -(11)/2+75=69.50